### v2

@@ 4.5 Post-process optimization
Anchor-based detection models typically rely on Non-Maximum Suppression (NMS) as a core post-processing operation to filter redundant detection boxes using an Intersection over Union (IoU) threshold. Traditional NMS performs robustly in sparsely distributed object scenarios but exhibits inherent limitations in dense deformed table cell detection: when adjacent cells exhibit highly overlapping bounding boxes due to geometric deformation, the IoU criterion mistakenly identifies them as the same instance, thereby suppressing correctly localized boxes with lower confidence scores. Although Soft-NMS [42] partially alleviates over-suppression through a confidence decay mechanism, severe overlaps caused by complex target shapes still lead to inflated IoU values, failing to address the root cause. To resolve this, inspired by the mask competition strategy in SOLOv2 [43], we abandon the bounding-box suppression paradigm and introduce Mask-Driven Non-Maximum Suppression, which employs pixel-level Mask_IoU for redundancy elimination. The metric is defined as:
@@ 5.2 Evaluation metrics
To comprehensively evaluate the performance of instance segmentation models, we adopt general metrics including mean Average Precision (mAP), model parameter count, and GFLOPs as quantitative criteria. mAP measures the model’s segmentation and classification capabilities by calculating average precision across multiple confidence thresholds. Specifically, precision P@t and recall R@t at a given confidence threshold t are first computed using Equation 18:
@@
where FP@t denotes false positives at confidence threshold t, and FN@t represents false negatives at the same threshold.
Average Precision (AP) is defined as the area under the precision-recall (PR) curve. The per-class average precision AP@t at confidence threshold t is calculated via Equation 19:
@@
Finally, the mean Average Precision mAP@t is obtained by averaging the AP@t values across all classes at confidence threshold t using Equation 20. Different confidence thresholds lead to distinct metrics: mAP@50 uses a confidence threshold of 0.5, reflecting baseline performance under lenient localization requirements, while mAP@50:95 rigorously evaluates the model’s robustness to object boundaries by averaging results across confidence thresholds from 0.5 to 0.95 with a step size of 0.05.
$\text{mAP}@t=\frac{1}{N_{\text{classes}}}\sum_{i=1}^{N_{\text{classes}}}\text{AP}_{i}@t$ (20)
The number of parameters (Parameters) reflects the model’s complexity and storage requirements. Excessive parameters may lead to overfitting and deployment challenges, while insufficient parameters can limit feature representation capacity. GFLOPs quantifies the computational demand per inference in billions of floating-point operations, measuring computational efficiency. Models with high GFLOPs rely on high-performance GPUs and struggle to meet real-time requirements, whereas low-GFLOPs designs are suitable for real-time video processing but require architectural optimization to balance accuracy trade-offs. All GFLOPs values in this study are reported under the input size of $640\times 640$.
@@ 5.3 Quantitative result
The experimental results on the DWTAL-s dataset, as shown in Table 1, demonstrate that the OG-HFYOLO model achieves a mAP@50:95 of 74.23%, outperforming mainstream two-stage instance segmentation models such as Mask R-CNN (62.5%) and Cascade Mask R-CNN (62.1%). It surpasses classical YOLOv8 (57.5%) and the more advanced YOLOv11 (57.8%) by 16.73% and 16.43%, respectively, and exceeds the highest-precision YOLOv5 variant (71.96%) by 2.27%. Additionally, the proposed model outperforms the Transformer-based Mask2Former (63.3%) by 10.93%. Regarding model parameter count, the introduction of a heterogeneous kernel cross-fusion architecture with relatively large convolutional kernels slightly increases the parameter size, though this remains manageable under current hardware storage constraints. In terms of inference speed, YOLACT has seriously lost model accuracy in the pursuit of extreme speed and model size, while the OG-HFYOLO model balances accuracy and speed, retaining the speed that a single-stage model should have. It surpasses both mainstream two-stage models and some single-stage models like SOLOv2 and Mask2Former in speed.
Table 1: The different model performance on DWTAL-s
@@
Table 2: The different model performance on DWTAL-l
@@ 5.4 Qualitative analysis
Figure 9: Our model result in real-world and other dataset
@@ 5.5.1 Impact of Anchor Mechanism
With the advancement of deep learning, detection-oriented models have progressively transitioned from anchor-based to anchor-free [52] mechanisms to pursue higher speed and smaller parameter footprints. Starting from YOLOv5, models began reducing reliance on predefined anchors, and by YOLOv8, anchor-free mechanisms became the standard. Comparative experiments clearly demonstrate that YOLOv8l-seg and YOLOv11l-seg, which adopt anchor-free mechanisms, exhibit significantly lower accuracy than their YOLOv5l-seg counterparts within the same YOLO series. To validate the superiority of anchor-based mechanisms for the current task, ablation studies on anchor mechanisms were conducted. As shown in Table 3, across both derived datasets, all metrics for both YOLOv5 and OG-HFYOLO models achieve over 10% significant improvement when using anchor-based mechanisms, highlighting anchor-based effectiveness in this research context.
Table 3: Impact of Anchor Mechanism
@@ 5.5.2 Effect of the proposed method
Table 4: Effect of the proposed method
@@ 5.6 Effect of module in the backbone
To clarify why the proposed model does not adopt the more advanced backbone architecture from YOLOv11, which introduces the flexible C3k2 module [51] for feature extraction and integrates a C2PSA attention mechanism [51] at the final stage to consolidate backbone features, ablation studies were conducted across different backbone configurations under identical experimental settings. As demonstrated in Table 5, the baseline C3 architecture achieves optimal performance on both datasets, outperforming variants incorporating YOLOv11’s enhanced components. The experimental results demonstrate that optimal results can be achieved by using the C3 module directly.
Table 5: Effect of different backbone
@@ 6 Conclusion
To obtain fine-grained spatial coordinates of cells in deformed tables, we introduce the DWTAL dataset. Addressing the challenges of dense object distributions and extreme scale variations within this dataset, the OG-HFYOLO model is proposed for accurate instance segmentation of tabular cells. The model incorporates several key innovations: the Gradient Orientation-aware Extractor (GOE) enhances contour perception for densely packed objects, the Heterogeneous Kernel Cross-Fusion (HKCF) and scale-aware Loss mitigate the challenges brought by severe scale variation, and a Mask-based Non-Maximum Suppression (MASK-NMS) is employed to prevent erroneous suppression caused by overlapping bounding boxes.
This work leverages instance segmentation to acquire precise spatial coordinates of cells in deformed tables, which is a mid-to-upstream task in table structure recognition. Downstream tasks involving logical coordinate processing can adopt the operational framework of LGPMA, optimizing the workflow through systematic integration of Computer Graphics principles and geometric topology theory. Furthermore, the challenges encountered in the derived dataset, such as dense object arrangements and scale diversity, are prevalent in domains like medical cell segmentation and remote sensing image analysis. Consequently, the proposed framework offers constructive insights for addressing analogous challenges across these fields.

### v3

@@ 4.5 Post-process optimization
Anchor-based detection models typically rely on Non-Maximum Suppression (NMS) as a core post-processing operation to filter redundant detection boxes using an Intersection over Union (IoU) threshold. Traditional NMS performs robustly in sparsely distributed object scenarios but exhibits inherent limitations in dense deformed table cell detection: when adjacent cells exhibit highly overlapping bounding boxes due to geometric deformation, the IoU criterion mistakenly identifies them as the same instance, thereby suppressing correctly localized boxes with lower confidence scores. Although Soft-NMS [42] partially alleviates over-suppression through a confidence decay mechanism, severe overlaps caused by complex object shapes still lead to inflated IoU values, failing to address the root cause. To resolve this, inspired by the mask competition strategy in SOLOv2 [43], we abandon the bounding-box suppression paradigm and introduce Mask-Driven Non-Maximum Suppression, which employs pixel-level Mask_IoU for redundancy elimination. The metric is defined as:
@@ 5.2 Evaluation metrics
To comprehensively evaluate the performance of instance segmentation models, we adopt general metrics including mean Average Precision (mAP), model parameter count, and GFLOPs as quantitative criteria. mAP measures the model’s segmentation and classification capabilities by calculating average precision across multiple confidence thresholds. Specifically, precision P(t) and recall R(t) at a given confidence threshold t are first computed using Equation 19:
@@
where FP(t) denotes false positives at confidence threshold t, and FN(t) represents false negatives at the same threshold.
Average Precision (AP) is defined as the area under the precision-recall (PR) curve. The per-class average precision AP(t) at confidence threshold t is calculated via Equation 20:
@@
Finally, the mean Average Precision mAP(t) is obtained by averaging the AP(t) values across all classes at confidence threshold t using Equation 21. Different confidence thresholds lead to distinct metrics: mAP@50 uses a confidence threshold of 0.5, reflecting baseline performance under lenient localization requirements, while mAP@50:95 rigorously evaluates the model’s robustness to object boundaries by averaging results across confidence thresholds from 0.5 to 0.95 with a step size of 0.05.
$\text{mAP}(t)=\frac{1}{N_{\text{classes}}}\sum_{i=1}^{N_{\text{classes}}}\text{AP}_{i}(t)$ (21)
The number of parameters reflects the model’s complexity and storage requirements. Excessive parameters may lead to overfitting and deployment challenges, while insufficient parameters can limit feature representation capacity. GFLOPs quantifies the computational demand per inference in billions of floating-point operations, measuring computational efficiency. Models with high GFLOPs rely on high-performance GPUs and struggle to meet real-time requirements, whereas low-GFLOPs designs are suitable for real-time video processing but require architectural optimization to balance accuracy trade-offs. All GFLOPs values in this study are reported under the input size of $640\times 640$.
@@ 5.3 Quantitative result
The experimental results on the DWTAL-s dataset, as shown in Table 2, demonstrate that the proposed OG-HFYOLO model achieves a mAP@50:95 of 74.23%, outperforming mainstream two-stage instance segmentation models such as Mask R-CNN (62.5%) and Cascade Mask R-CNN (62.1%). It surpasses classical YOLOv8 (57.5%) and the more advanced YOLOv11 (57.8%) by 16.73% and 16.43%, respectively, and exceeds the highest-precision YOLOv5 variant (71.96%) by 2.27%. Additionally, the proposed model outperforms the Transformer-based Mask2Former (63.3%) by 10.93%. Regarding model parameter count, the introduction of a heterogeneous kernel cross-fusion architecture with relatively large convolutional kernels slightly increases the parameter size, though this remains manageable under current hardware storage constraints. In terms of inference speed, YOLACT has seriously lost model accuracy in the pursuit of extreme speed and model size, while the OG-HFYOLO model balances accuracy and speed, retaining the speed that a single-stage model should have. It surpasses both mainstream two-stage models and some single-stage models like SOLOv2 and Mask2Former in speed.
Table 2: The different model performance on DWTAL-s. The best results are in bold.
@@
Table 3: The different model performance on DWTAL-l. The best results are in bold.
@@ 5.4 Qualitative analysis
Figure 9: Our model result in real-world and other dataset.
@@ 5.5.1 Impact of Anchor Mechanism
With the advancement of deep learning, detection-oriented models have progressively transitioned from anchor-based to anchor-free [52] mechanism to pursue higher speed and smaller parameter footprints. Starting from YOLOv5, models began reducing reliance on predefined anchors, and by YOLOv8, anchor-free mechanism became the standard. Comparative experiments clearly demonstrate that YOLOv8l-seg and YOLOv11l-seg, which adopt anchor-free mechanism, exhibit significantly lower accuracy than their YOLOv5l-seg counterparts within the same YOLO series. To validate the superiority of anchor-based mechanism for the current task, ablation studies on anchor mechanisms were conducted. As shown in Table 4, across both derived datasets, all metrics for both YOLOv5 and OG-HFYOLO models achieve over 10% significant improvement when using anchor-based mechanism, highlighting anchor-based effectiveness in this research context.
Table 4: Impact of Anchor Mechanism. The best results are in bold.
@@ 5.5.2 Effect of the proposed method
Table 5: Effect of the proposed method. The best results are in bold.
@@ 5.6 Effect of module in the backbone
To clarify why the proposed model does not adopt the more advanced backbone architecture from YOLOv11, which introduces the flexible C3k2 module [51] for feature extraction and integrates a C2PSA attention mechanism [51] at the final stage to consolidate backbone features, ablation studies were conducted across different backbone configurations under identical experimental settings (including the same anchor mechanism, training strategy, and proposed modules). As demonstrated in Table 6, the baseline C3 architecture achieves optimal performance on both datasets, outperforming variants incorporating YOLOv11’s enhanced components. The experimental results demonstrate that optimal results can be achieved by using the C3 module directly.
Table 6: Effect of different backbone. The best results are in bold.
@@ 6 Conclusion
To address spatial coordinate localization in deformed table structures, we propose a novel instance segmentation framework. This approach establishes a novel pixel-level parsing paradigm that significantly outperforms traditional detection-based localization methods. Compared with traditional detection methods, we establish a new localization paradigm grounded in pixel-level parsing. To facilitate this research direction, we construct the DWTAL dataset specifically designed for deformed table cell localization tasks. Its fine-grained spatial coordinate annotations fill the critical benchmark data gap in this research domain. To address the challenges of dense object distributions and extreme scale variations in DWTAL dataset, we further propose the OG-HFYOLO model for accurate instance segmentation of deformed table cells. The model incorporates several key innovations: the Gradient Orientation-aware Extractor (GOE) enhances contour perception for densely packed objects; the Heterogeneous Kernel Cross-Fusion (HKCF) and scale-aware Loss address challenges posed by severe scale variation; and Mask-based Non-Maximum Suppression (MASK-NMS) prevents erroneous suppression caused by overlapping bounding boxes.
This work leverages instance segmentation to acquire precise spatial coordinates of cells in deformed tables, a mid-to-upstream task in table structure recognition. Downstream tasks involving logical coordinate processing can adopt the operational framework of LGPMA, optimizing the workflow through systematic integration of Computer Graphics principles and geometric topology theory. Furthermore, challenges encountered in the derived dataset, such as dense object arrangements and scale diversity, are common in fields like medical cell segmentation and remote sensing image analysis. Consequently, the proposed methods provide constructive insights for addressing similar challenges in these fields.
Acknowledgments
This work is supported by the Natural Science Foundation of China (No. 62366034).